solve for x.(x-1)^3+(x-2)^3+(x-3)^3+(x-4)^3+(x-5)^3 in short teckniq.

Grade:12

3 Answers

Akash Kumar Dutta
10 years ago

where is = sign for solving?

Madhukar Thalore
10 years ago

Thjis is an example of sum of cubes.just use the formulae for the sum of cubes of first n natural numbers.

ALL THE BEST.

jaymin prajapati
10 years ago

i got answer. Here A.M of {x-1, x-5} =A.M of {x-2, x-4} =x-3, put x-3=y then,
x-1= y+2, x-2=y+1, x-4=y-1, x-5=y-2.



The equation becomes,